# A Posteriori Single- and Multi-Goal Error Control and Adaptivity for Partial Differential Equations

### v1

@@ 5.3 Error Localization
$J_{\mathfrak{E}}(u)-J_{\mathfrak{E}}(\tilde{u})\approx\eta^{(2)}=\sum_{i=1}^{M}\frac{1}{2}\rho(\tilde{u})((z_{h}^{(2)}-\tilde{z})\chi_{i})+\frac{1}{2}\rho^{*}(\tilde{u},\tilde{z})((u_{h}^{(2)}-\tilde{u})\chi_{i})+\rho(\tilde{u})(\tilde{z})$
@@ 6.1.4 Goal Functional (Quantity of Interest) and Adjoint Problem
and was obtained on a globally refined super-mesh.
@@ 6.4.2 Space-Time Finite Element Discretization and Non-Linear Solver
We implemented the space-time finite element method using the finite element library MFEM [7]. We use a $P_{1}$ conforming ansatz space for the 2D+1 ($d=2$) and 3D+1 ($d=3$) case, i.e. we use a simplicial decomposition $\mathcal{T}_{h}$ of the $(d+1)$-dimensional space-time cylinder $Q\subset\mathbb{R}^{d+1}$.
@@ 7 Conclusions and Outlook
In this work, we reviewed single- and multigoal-oriented error control and adaptivity. Our notation was kept in general terms such that stationary and non-stationary (space-time) situations are covered. First, we explained in detail single goal-oriented error estimation with the help of the dual-weighted residual method. The error estimators cover both discretization and non-linear iteration errors. Therein, some new theoretical results were presented as well. Prior efficiency and reliability results only require one saturation assumption, rather a strengthened condition. Next, we considered non-standard discretizations such as stabilization terms and non-consistencies (e.g., classical finite difference based time-stepping schemes) or non-conformal methods (e.g., finite differences or neural network approximations) of numerical schemes in the frame of goal-oriented error estimation. Third, we concentrated on multigoal-oriented error estimation in which we have put a lot of efforts in the last eight years due to advances and increasing interest of solving multiphysics partial differential equations and coupled variational inequality systems. Besides theoretical results, we provided several adaptive algorithms for single and multiple goal functional evaluations. In order to substantiate our developments, four numerical examples were designed and computationally analyzed. In the first example, namely Poisson’s problem, the implementation is fully open-source and follows the classical structure of deal.II tutorial steps. Therefore, this example has a flavor of educational purpose. In the last example, recent work on the space-time PU-DWR method is further extended to multiple-goal functionals and singularities in the solution. Ongoing and future work is concerned with the extension to cover more terms in the error estimators that besides discretization and non-linear iteration errors, as well linear iteration errors, model errors, and model order reduction techniques are covered. In conjunction with high-performance parallel computing these yield important components to continue to solve efficiently and with desired accuracies multiphysics problems with physics-based discretizations and fast physics-based numerical solvers.

### v2

@@ 5.3 Error Localization
$J_{\mathfrak{E}}(u)-J_{\mathfrak{E}}(\tilde{u})\approx\eta^{(2)}=\sum_{i=1}^{M}\frac{1}{2}\rho(\tilde{u})((z_{h}^{(2)}-\tilde{z})\chi_{i})+\frac{1}{2}\rho^{*}(\tilde{u},\tilde{z})((u_{h}^{(2)}-\tilde{u})\chi_{i})+\rho(\tilde{u})(\tilde{z}),$
@@ 6.1.4 Goal Functional (Quantity of Interest) and Adjoint Problem
and was obtained on a globally refined super-mesh. In this simple example, with the ansatz $u(x,y)=\sin(\pi x)\sin(\pi y)$, we could have constructed a corresponding right hand side $f$ and with that we would have obtained an analytical reference value. Computing reference values on super-refined meshes is usually the method when no analytical or manufactured solutions are available.
@@ 6.4.2 Space-Time Finite Element Discretization and Non-Linear Solver
We implemented the space-time finite element method using the finite element library MFEM [7]. We use a $P_{1}$ conforming ansatz space for the 2D+1 ($d=2$) and 3D+1 ($d=3$) case, i.e. we use a simplicial decomposition $\mathcal{T}_{h}$ of the $(d+1)$-dimensional space-time cylinder $Q\subset\mathbb{R}^{d+1}$. For $d=2$, i.e. $Q\subset\mathbb{R}^{3}$, we use the bisection algorithm by Arnold, Mukherjee and Pouly [11], while for $d=3$, i.e. $Q\subset\mathbb{R}^{4}$, we use the newest vertex bisection from Stevenson’s paper [212]. Both algorithms ensure the shape regularity of the generated sequence of meshes.
@@ 7 Conclusions and Outlook
In this chapter, we reviewed single- and multigoal-oriented error control and adaptivity. Our notation was kept in general terms such that stationary and non-stationary (space-time) situations are covered. First, we explained in detail single goal-oriented error estimation with the help of the dual-weighted residual method. The error estimators cover both discretization and non-linear iteration errors. Therein, some new theoretical results were presented as well. Prior efficiency and reliability results only require one saturation assumption, rather a strengthened condition. Next, we considered non-standard discretizations such as stabilization terms and non-consistencies (e.g., classical finite difference based time-stepping schemes) or non-conformal methods (e.g., finite differences or neural network approximations) of numerical schemes in the frame of goal-oriented error estimation. Third, we concentrated on multigoal-oriented error estimation in which we have put a lot of efforts in the last eight years due to advances and increasing interest of solving multiphysics partial differential equations and coupled variational inequality systems. Besides theoretical results, we provided several adaptive algorithms for single and multiple goal functional evaluations. In order to substantiate our developments, four numerical examples were designed and computationally analyzed. In the first example, namely Poisson’s problem, the implementation is fully open-source and follows the classical structure of deal.II tutorial steps. Therefore, this example has a flavor of educational purpose. In the last example, recent work on the space-time PU-DWR method is further extended to multiple-goal functionals and singularities in the solution.
Ongoing and future work is concerned with the extension to cover more terms in the error estimators that besides balancing discretization and non-linear iteration errors, as well linear iteration errors, model errors, and model order reduction techniques are covered. Such error estimators could be theoretically investigated again with the help of saturation assumptions. The development and the analysis of fast iterative solvers for the linear and non-linear systems of finite element equations arising in the single- and multi-goal DWR AFEM considered in this chapter are also interesting topics of further research work, in particular, for the space-time AFEM presented in Subsection 6.4. It is very natural to use a nested iteration framework that reuses the generated sequence of adaptive meshes for the solvers like in the full linear and non-linear multigrid solvers; see, e.g., [124], or [53] for the full approximation scheme (FAS). Moreover, one should think about smart techniques between the use of expensive enriched spaces based on refined meshes or high-order elements and interpolation that keep the quality of enriched spaces with the efficiency of the interpolation technique. In conjunction with high-performance parallel computing these yield important components to continue to solve efficiently and with desired accuracies multiphysics problems with physics-based discretizations and fast physics-based numerical solvers. Two such applications are space-time fluid-structure interaction and space-time dynamic phase-field fracture problems. Partial results are available to date, but full space-time adaptivity not yet. Both can serve as full order models within model order reduction, where the reduced order models are designed themselves with the help of goal-oriented error estimation. Such by far cheaper reduced models enhance significantly the efficiency within optimization loops such as optimal control, inverse design and parameter estimation.